OpenStudy (anonymous):

(a^-3 b^2/a^-6 b ^10)^-4

OpenStudy (anonymous):

do you know what a negative exponent requires you to do to this.

OpenStudy (haseeb96):

\[[\frac{ a ^{-3 }b ^{2} }{ a ^{-6} b ^{10} }]^{-4}\] \[[\left( a ^{-3+6}b ^{2-10} \right)]^{-4}\] \[[\left( a ^{3}b ^{-8} \right)]^{-4}\] \[a ^{-12}b^{32}\] is the correct answer

OpenStudy (anonymous):

Thank you you are my life saver